# Talk2Radar: Bridging Natural Language with 4D mmWave Radar for 3D Referring Expression Comprehension

### v1

@@ IV-C Deformable FPN
To enhance representation of text-conditioned three-stage point cloud features , we first exert the efficient deformable convolutions [52] on three maps to perform adaptive sparse sampling and modeling of key features. For the current element in the one-stage radar map, the deformable convolution can be formulated as Eq. (7):
@@
Then and are upsampled to the same size as by transposed convolution. Finally, three-scale feature maps are concatenated to obtain high-resolution aggregated point cloud features with multi-receptive fields.
@@ IV-D Training Objectives
The center-based detection head first produces class-wise heatmaps to predict the center location of the detected objects for each class. Afterward, the following properties of each object are supervised following CenterPoint [48]: the sub-voxel location refinement, the height-above-ground, the 3D size, and the orientation angle. The proposed T-RadarNet is trained with the following loss:
where is the classification loss that supervises the heatmap quality of the center-based detection head using a focal loss [54], and BB = which indicates the smooth-L1 loss supervises the regression of the box center (for slight modification based on heatmap peak guidance), dimensions, and orientation. is the weight to balance the two components of the loss, which is set to 0.25 by default.
@@ V-A Implementation Settings
Model Settings. Firstly, for our T-RadarNet, we set the token length as 30 upon ALBERT [55]. in the pillar encoder is set as 10 and 32 for radar and LiDAR, respectively while is 64. We select PC-based detectors with various paradigms for comparison and analysis, including PointPillars (pillar-based) [42], CenterPoint (anchor-free) [48], SECOND (voxel-based) [46] and CenterFormer (transformer-based) [49]. Further, we also include dedicated 3D PC grounding model MSSG [31] and AFMNet [41]. Secondly, fusion methods including inductive bias-based HDP [56], attention-based MHCA [28] and our GGF. Thirdly, SECONDFPN [46], ASPP [43] and CSP-FPN [57] are included for comparison with Deformable FPN.
Dataset Settings. We mainly focus on objects of three primary categories, i.e., Car, Cyclist, and Pedestrian, in Talk2Radar while we also train and test models on all categories (in supplementary materials). We train and test models on both 4D radar and LiDAR data. Further, to validate the generalization of our T-RadarNet, we also train and test it on Talk2Car [36], where it provides LiDAR PC and each sample contains one textual prompt for a unique object.
Training and Evaluation Settings. For Talk2Radar, all models are trained on four RTX A4000 with a batch size of 4 for 80 epochs. The initial learning rate is set 1e-3 with a cosine scheduler. We choose AdamW for optimization with a weight decay of 5e-4. For Talk2Car, we set the initial learning rate as 1e-2 while other settings are kept the same with Talk2Radar.
@@ V-B Quantitative Results
Overall Performances. As Table III shows, on the whole, T-RadarNet outperforms other models in most aspects whatever on 4D radar or LiDAR. For other models, we find that models based on pillar encoding perform better than those based on voxels, and models based on self-attention do not significantly outperform those based on convolution. Regarding fusion, models using MHCA perform worse than those using gating and dot-product operations. This phenomenon is particularly evident with five-frame radar data (Radar5).
Performances on Various Prompts. As Table IV shows, when we compare the performance of T-RadarNet on prompts with queries of types on motion, depth and velocity, the radar-based model outperforms LiDAR-based on prompts regarding motion and velocity, which is more obvious on velocity-based prompts. Besides, as LiDAR can reason the object motion upon PC object appearances, the gap between radar and LiDAR on motion query is not large. For the depth-related prompts, LiDAR obtains the advantage to some extent.
Precision by Piece-wise Depths. Table V presents the mAP of predicted objects by segmented depths upon 5-frame radar. Small-object grounding is doubtless a challenge.
Generalization Performances. To validate the generalization of T-RadarNet, we train and test on Talk2Car, which is a REC benchmark providing LiDAR data for object grounding. As Table VII shows, our T-RadarNet outperforms the baseline and the other two models, which proves the effectiveness and generalization of T-RadarNet for PC-based 3D REC.
@@
| Fusion Methods | ||||||
| HDP | 20.07 | 6.52 | 14.86 | 39.53 | 8.62 | 15.85 |
| MHCA | 7.79 | 5.02 | 5.36 | 13.69 | 5.81 | 9.35 |
| GGF | 24.68 | 9.71 | 15.74 | 42.58 | 10.13 | 17.82 |
| Fusion Locations | ||||||
| neck-head | 17.53 | 8.10 | 12.06 | 38.56 | 5.03 | 14.59 |
| backbone-neck | 24.68 | 9.71 | 15.74 | 42.58 | 10.13 | 17.82 |
| Models | BEV AP | 3D AP | ||
| AP | AP | AP | AP | |
| Talk2Car [36] | 30.6 | 24.4 | 27.9 | 19.1 |
| MSSG [31] | 27.8 | 26.1 | 31.9 | 20.3 |
| AFMNet [41] | 45.3 | 33.1 | 41.9 | 20.7 |
| T-RadarNet | 52.8 | 39.9 | 47.2 | 30.5 |
@@ V-C Ablation Experiments
Table VI presents the ablation results of T-RadarNet.
GGF. When we replace the MRConv4D (GConv) with normal convolution, mAP drops remarkably, which is the same circumstance with pooling for textual semantics abstraction.
Necks. Our Deformable FPN outperforms other necks overall, indicating the preeminent modeling of point clouds.
Fusion Methods. Our proposed GGF presents a better result than other fusion methods, implying a more efficient approach to aligning and embedding textual semantics to point clouds. Although MHCA can model global cross-modal similarity, its performance is sub-optimal when dealing with irregular point clouds containing both targets and ghosting in radar data.
@@ V-D Visualization and Discussion
Fig. 5 presents the prediction results by T-RadarNet based on the 5-frame accumulated radar PC, including visualization on the image plane and BEV view in the PC context. For the correct prediction results in the first row, we can see our T-RadarNet possesses the capacity to understand the textual prompts with different lengths and localize the correct objects among complex point cloud contexts, including single and multiple referent objects. However, there are still some problems with this task, in the second row, we can see the serious false positive cases in the prediction bounding boxes, implying that efficient modeling of radar PC features and adaptive filtering of non-object clutter are still a challenge.
@@ VI Conclusion
This paper proposes a novel task: 4D mmWave radar-based 3D referring expression comprehension (visual grounding), aimed at embodied intelligence and interactive perception with full weather operational, low-cost sensor. The first dataset on such task, Talk2Radar, has been presented by fully exploiting object detection characteristics of 4D mmWave radar. Talk2Radar includes abundant textual prompts and object distributions, as well as LiDAR. Through extensive experiments, the corresponding benchmark has been provided by proposing an efficient radar-based 3D REC model, T-RadarNet. In T-RadarNet, we design an effective module, Gated Graph Fusion, for the alignment and fusion of textual and 4D radar point cloud features. Additionally, Deformable FPN is proposed to adequately model irregular and sparse point cloud features. Having all the aforementioned contributions, we aim to advance the interactive perception capabilities of 4D radar for environmental understanding in autonomous driving.

### v3

@@ IV-C Deformable FPN
To enhance representation of text-conditioned three-stage PC features , we first exert the efficient deformable convolutions [62] on three maps to perform adaptive sparse sampling and modeling of key features. For the current element in the one-stage radar map, the deformable convolution can be formulated as
@@
Then and are upsampled to the same size as by transposed convolution. Finally, three-scale feature maps are concatenated to obtain high-resolution aggregated PC features with multi-receptive fields.
@@ IV-D Training Objectives
The center-based detection head first produces class-wise heatmaps to predict the center location of the detected objects for each class. Afterward, the following properties of each object are supervised following CenterPoint [58]: the sub-voxel location refinement, the height-above-ground, the 3D size, and the orientation angle. The proposed T-RadarNet is trained with the following loss
where is the classification loss supervising the heatmap quality of the center-based detection head using a focal loss [68]; indicates the smooth- loss supervising the regression of the box center (for slight modification based on heatmap peak guidance), dimensions, and orientation; and is the weight to balance the two components of the loss, which is set to 0.25 by default.
@@ V-A Implementation Settings
Model Settings: For T-RadarNet, we set the token length as 30 upon ALBERT [64] and other text encoders for comparison. in the pillar encoder is set as 10 and 32 for radar and LiDAR, respectively, while is 64. For comparison, we firstly select PC-based detectors with various paradigms including PointPillars (pillar-based) [52], CenterPoint (voxel and anchor-free) [58], and CenterFormer (transformer-based) [59]. Secondly, fusion methods including inductive bias-based HDP [69], attention-based MHCA [32] are implemented to compare with our GGF. Thirdly, different necks, i.e., SecondFPN [56], ASPP [53] and CSP-FPN [70] are included for comparison with proposed Deformable FPN. Lastly, we also compare proposed T-RadarNet with dedicated SOTA 3D PC grounding model MSSG [35], AFMNet [46] and EDA [66].
Dataset Settings: We train and test models on Talk2Radar dataset with three type of objects, i.e., Car, Cyclist, and Pedestrian, using both 4D radar and LiDAR. Moreover, to validate the generalization of T-RadarNet, we also train and test it on Talk2Car dataset [41], which was built upon nuScenes [42] by providing textual prompt for LiDAR PC.
Training and Evaluation Settings: For Talk2Radar, all models are trained on four RTX A4000 with a batch size of 4 for 80 epochs. The initial learning rate is set 1e-3 with a cosine scheduler. We choose AdamW for optimization with a weight decay of 5e-4. For Talk2Car, we set the initial learning rate as 1e-2 while other settings are kept the same with Talk2Radar.
@@ V-B Quantitative Results
Overall Performances: As Table III shows, on the whole, T-RadarNet outperforms other models in most aspects whatever on 4D radar or LiDAR. For other models, we find that the models based on pillar encoding perform better than those based on voxels, and the models based on self-attention do not significantly outperform those based on convolution. Regarding fusion, models using MHCA perform worse than those using gating and dot-product operations. This phenomenon is particularly evident with five-frame radar data (Radar5). Performances on Various Prompts: As Table IV shows, when we compare the performance of T-RadarNet on prompts with queries of types on motion, depth and velocity, the radar-based model outperforms LiDAR-based on prompts regarding motion and velocity, which is more obvious on velocity-based prompts. Besides, as LiDAR can reason the object motion upon PC object appearances, the gap between radar and LiDAR on motion query is not large. For the depth-related prompts, LiDAR obtains the advantage to some extent. Precision by Piece-wise Depths: Table V presents the mAP of predicted objects by segmented depths upon 5-frame radar. Small-object grounding is doubtless a challenge. Generalization Performances: Table VI shows the generalization evaluation results of T-RadarNet on Talk2Car for 3D REC, our T-RadarNet outperforms the baseline and the other three models, which proves the effectiveness and generalization of T-RadarNet for PC-based 3D REC.
@@
| Text Encoders | ||||||
| Bi-GRU [65] | 18.74 | 7.57 | 10.89 | 24.60 | 7.98 | 11.50 |
| RoBERTa [67] | 24.22 | 9.53 | 16.12 | 42.62 | 10.03 | 16.66 |
| ALBERT | 24.68 | 9.71 | 15.74 | 42.58 | 10.13 | 17.82 |
@@ V-C Ablation Experiments
Table VII presents the ablation results of T-RadarNet. GGF: When we replace the MRConv4D (GConv) with normal convolution, mAP drops remarkably, which is the same circumstance with pooling for textual semantics abstraction. Fusion Methods: Our proposed GGF presents a better result than other fusion methods, implying a more efficient approach to aligning and embedding textual semantics to PCs. Although MHCA can model global cross-modal similarity, its performance is sub-optimal when dealing with irregular PCs containing both objects and ghosting in radar data. Necks: Deformable FPN outperforms all other necks, indicating its preeminent capability of representing PCs. Text Encoders: Pretrained transformer-based text encoders outperform RNN-based text encoders remarkably, indicating the significance of attention weighting on text encoding for the following fusion.
@@ V-D Visualization and Discussion
Fig. 5 presents the prediction results by T-RadarNet based on the 5-frame accumulated radar PC, including visualization on the image plane and BEV view in the PC context. For the correct prediction results in the first row, we can see our T-RadarNet can understand textual prompts and localize the correct objects among radar PC contexts, including single and multiple referent objects. However, there are still some problems with this task, in the second row, we can see the serious false positive cases in the predicted bounding boxes, implying that efficient representation of radar PC features and adaptive filtering of non-object clutter are still a challenge.
@@ VI Conclusion
This paper proposes a novel task: 4D mmWave radar-based 3D referring expression comprehension (3D visual grounding), aimed at enhancing embodied intelligence and interactive perception with an all-weather, low-cost sensor. We introduce the first dataset for this task, Talk2Radar, which fully leverages the object detection capabilities of 4D mmWave radar. Talk2Radar includes a rich collection of textual prompts and object distributions, alongside LiDAR data. Through extensive experiments, we establish a corresponding benchmark by proposing an efficient radar-based 3D REC model, T-RadarNet. Within T-RadarNet, we design an effective module, Gated Graph Fusion, for the alignment and fusion of textual and 4D radar point cloud features. Additionally, we propose Deformable FPN to adequately model irregular and sparse point cloud features. With these contributions, we aim to advance the interactive perception capabilities of 4D radar for environmental understanding in autonomous driving.